Two concentric current loops lie in the same plane. The smaller loop has a radius of 3.0 cm and a current of 12 A. The bigger loop has a current of 20 A. The magnetic field at the center of the loops is found to be zero.
What is the radius of the bigger loop?
Answer #1
Concepts and reason

The concept used to solve this problem is Biot-Savart’s law.

At first, use the concept of Biot-savart’s law to derive the expression for the magnetic field at the center of the circular coil carrying current and then, use this expression to determine the radius of the bigger loop.

Fundamentals

Biot-Savart’s law:

This law is used to find the magnetic field at a point due to a current carrying element.

According to this law, the magnitude of the magnetic field induction at a point BB , which is at a distance rr from the central axis of the wire is:

• Directly proportional to the current II in the element, dBIdB \propto I

• Directly proportional to the small element dldl of the length of the conductor ll , dBdldB \propto dl

• Directly proportional to the sine of the angle θ\theta between the line joining the element dldl and the point BB , dBsinθdB \propto \sin \theta

• Inversely proportional to the square of the distance rr from the current carrying element dldl to the point BB , dB1r2dB \propto \frac{1}{{{r^2}}}

Here, dBdB is the magnitude of the magnetic field at point BB .

The following diagram explains the concept of Biot-savart’s law.

dhe

Thus, dBIdlsinθr2dB \propto \frac{{Idl\sin \theta }}{{{r^2}}}

Replace the proportionality sign with a proportionality constant.

dB=μ04πIdlsinθr2dB = \frac{{{\mu _0}}}{{4\pi }}\frac{{Idl\sin \theta }}{{{r^2}}}

Here, μ04π\frac{{{\mu _0}}}{{4\pi }} is the proportionality constant and μ0{\mu _0} is the permeability of the free space.

According to Biot-Savart’s law, the magnetic field at a point BB due to a current carrying element is given as:

dB=μ04πIdlsinθr2dB = \frac{{{\mu _0}}}{{4\pi }}\frac{{Idl\sin \theta }}{{{r^2}}} ...... (1)

The magnetic field at the center of the circular coil carrying current can be calculated with the help of the following diagram:

F
0

The angle between dl\overrightarrow {dl} and r\overrightarrow r is 90{90^ \circ } .

Substitute 90{90^ \circ } for θ\theta in the above equation,

dB=μ04πIdlsin90r2dB = \frac{{{\mu _0}}}{{4\pi }}\frac{{Idl\sin {{90}^ \circ }}}{{{r^2}}}

dB=μ04πIdlr2dB = \frac{{{\mu _0}}}{{4\pi }}\frac{{Idl}}{{{r^2}}}

The total magnetic field at a point OO in the circular coil can be obtained by integrating the above equation.

B=μ04πIr2dlB = \frac{{{\mu _0}}}{{4\pi }}\frac{I}{{{r^2}}}\int {dl}

The total length of the circular coil is 2πr2\pi r .

Substitute 2πr2\pi r for dl\int {dl} in the above equation.

B=μ0I2rB = \frac{{{\mu _0}I}}{{2r}}

This is the required magnetic field at the center of the circular coil carrying current.

The magnetic field at the center of the circular coil carrying current is given by:

B=μ0I2rB = \frac{{{\mu _0}I}}{{2r}}

The magnetic field of the two concentric loops is the same.

Thus, B1=B2{B_1} = {B_2}

Here, B1{B_1} is the magnetic field of the smaller loop and B2{B_2} is the magnetic field of the bigger loop.

Substitute μ0I12r1\frac{{{\mu _0}{I_1}}}{{2{r_1}}} for B1{B_1} and μ0I22r2\frac{{{\mu _0}{I_2}}}{{2{r_2}}} for B2{B_2} .

μ0I12r1=μ0I22r2I1r1=I2r2\begin{array}{l}\\\frac{{{\mu _0}{I_1}}}{{2{r_1}}} = \frac{{{\mu _0}{I_2}}}{{2{r_2}}}\\\\\frac{{{I_1}}}{{{r_1}}} = \frac{{{I_2}}}{{{r_2}}}\\\end{array}

Here, I1{I_1} is the current flowing through the smaller loop, I2{I_2} is the current flowing through the larger loop, r1{r_1} is the radius of the smaller loop and r2{r_2} is the radius of the larger loop.

Rearrange the above equation for r2{r_2} .

r2=I2r1I1{r_2} = \frac{{{I_2} \cdot {r_1}}}{{{I_1}}}

Substitute 12A12{\rm{ A}} for I1{I_1} , 3cm3{\rm{ cm}} for r1{r_1} and 20A20{\rm{ A}} for r2{r_2} .

r2=(20A)(3cm)(12A)=5cm\begin{array}{c}\\{r_2} = \frac{{\left( {20{\rm{ A}}} \right) \cdot \left( {3{\rm{ cm}}} \right)}}{{\left( {12{\rm{ A}}} \right)}}\\\\ = 5{\rm{ cm}}\\\end{array}

This is the required radius of the bigger loop.

Ans:

The radius of the bigger loop is 5cm5{\bf{ cm}} .